Find the surface area of a cylinder with radius 10 cm and height 13 cm.
A. 722.2 cm^2
B. 1444.4 cm^2
C. 1632.8 cm^2
D. 1583.4 cm^2
I applied the surface area formula of a cylinder and received the answer as 1445.13. I was wondering if B was the answer not. Thanks for helping!

Answer by Alan3354(69443)

Find the surface area of a cylinder with radius 10 cm and height 13 cm.
----------
Lateral area = 2pi*r*h = 260pi
Ends' area = 2*pi*r^2 = 200pi
----------------
B is only one that is close.